Synovec Co. is growing quickly. Dividends are expected to grow at a rate of 25 percent for the next three years, with the growth rate falling off to a constant 6 percent thereafter. If the required return is 11 percent, and the company just paid a dividend of $1.15, what is the current share price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
1
Expert's answer
2020-05-12T10:49:28-0400

Curent price of share can be calculated using two-stage divdend discount model:

"P=\frac{1.15*1.25}{1.11}+\frac{1.15*1.25^2}{1.11^2}+\frac{1.15*1.25^3}{1.11^3}+\frac{1.15*1.25^3*1.06}{1.11^3*0.05}=39.21"